### v1

# PDNNet: PDN-Aware GNN-CNN Heterogeneous Network for Dynamic IR Drop Prediction

Abstract
IR drop on the power delivery network (PDN) is closely related to PDN’s configuration and cell current consumption. As the integrated circuit (IC) design is growing larger, dynamic IR drop simulation becomes computationally unaffordable and machine learning based IR drop prediction has been explored as a promising solution. Although CNN-based methods have been adapted to IR drop prediction task in several works, the shortcomings of overlooking PDN configuration is non-negligible. In this paper, we consider not only how to properly represent cell-PDN relation, but also how to model IR drop following its physical nature in the feature aggregation procedure. Thus, we propose a novel graph structure, PDNGraph, to unify the representations of the PDN structure and the fine-grained cell-PDN relation. We further propose a dual-branch heterogeneous network, PDNNet, incorporating two parallel GNN-CNN branches to favorably capture the above features during the learning process. Several key designs are presented to make the dynamic IR drop prediction highly effective and interpretable. We are the first work to apply graph structure to deep-learning based dynamic IR drop prediction method. Experiments show that PDNNet outperforms the state-of-the-art CNN-based methods by up to $39.3\%$ reduction in prediction error and achieves 545×545$\times$545 × speedup compared to the commercial tool, which demonstrates the superiority of our method.
Index Terms:
Machine learning, IR drop prediction, Voltage drop analysis, Power grid analysis.
Refer to caption Refer to caption
(a) PDN with sparse grid design.
Refer to caption Refer to caption
(b) PDN with dense grid design.
Figure 1: Example of IR drop maps with different PDNs under the same current load. The left part is PDN architecture. High metal (e.g., M8 and M7) and low metal (e.g., M1) are distinguished with different color. The corresponding IR drop map is in the right position. The color approaching Red denotes a higher IR drop value, and approaching Blue denotes lower.
I Introduction
[…]
[lines=2]Power delivery network is responsible for transmitting voltage and current from the chip’s IO pins to every transistor  [1, 2, 3, 4, 5]. It contains a hierarchical architecture with a set of power bumps and multi-redistribution layers from top to bottom. Due to the existence of parasitics, a voltage drop is induced when the supply current flows through PDN. This common but unpleasant phenomenon is named IR drop, which is determined by two key factors (current load patterns as well as PDN architecture). The Figure 1 (a)(b) shows two IR drop distribution maps in distinctive PDN architectures under the same current load pattern. Large IR drops can hurt chip performance, timing, and in the worst case, its functionality. Thus, IR drop analysis is a crucial step to verify the satisfaction and reliability of PDN before manufacturing. [6, 7, 8, 9, 10, 11] As simulation-based analysis is too expensive, especially for dynamic IR drop, machine learning based IR drop prediction have been achieved to accelerate design closure.
In literature, convolutional neural networks (CNNs) in machine learning have dominated this task for a long time. Current CNN-based methods regard dynamic IR drop prediction as an image generation process [12, 13, 14, 15, 16]. Mozaffari et al. [17] develop an on-chip power supply noise evaluation method, leveraging the CNN model to handle global and local features in prediction. Xie et al. [18] propose a max-CNN structure, called PowerNet, that encodes input power features to capture the peak voltage drop along the temporal axis. Chhabria et al. [19] propose MAVIREC, which formulates the prediction task as an image-to-image translation problem, utilizing a UNet architecture to percept dynamic current variation.
Although these methods achieve good performance, the prominent drawbacks can not be neglected. As Figure 1 shows, current load patterns and PDN architectures are two indispensable parts that jointly impact the final IR drop distribution. Conventional works focus on extracting the representation from current load patterns, while almost none of them devote attention to explore the PDN structures. The reason mainly lies in two-fold. Firstly, image-like input features cannot fully represent PDN structure and current flow. Secondly, as pure CNN-based models lack topology relation aggregation approaches, they are neither capable of perceiving fine-grained cell-PDN connection relation, nor modeling the current flowing behavior on PDN.
This paper aims to target IR drop prediction from a brand-new angle based on the physical nature of this problem. The main contributions of this work are summarized below:
TABLE I: Comparison between various dynamic IR drop analysis methods.
Name Method Model Cell-PDN Relation Generality Speed
Voltus [20] Simulation IR Drop Analysis Yes High Medium
[21] Prediction XGBoost No Medium Medium
PowerNet [18] Prediction CNN No Medium Medium
MAVIREC [19] Prediction CNN No High High
Our PDNNet Prediction GNN-CNN Heterogeneous Yes (use PDNGraph) High High
  • We propose a novel graph structure, PDNGraph, which comprehensively represents the fine-grained PDN structure and cell-PDN relation. To the best of our knowledge, we are the first to introduce graph structure into deep-learning based dynamic IR drop prediction.
  • We propose a novel dual-branch network, PDNNet, that concentrates on two core perspectives. The GNN branch tailors the aforementioned PDNGraph from the perspective of PDN awareness. The parallel CNN branch focuses on capturing dynamic IR drop variation along the temporal axis.
  • We conduct extensive experiments on a large-scale dataset. PDNNet outperforms the state-of-the-art CNN-based methods by up to 39.3% reduction in prediction error and achieves 545$\times$ speedup compared to the commercial tool, which demonstrates the superiority of our proposed method.
The rest of the paper is organized as follows. The IR drop analysis procedure, problem formulation and method comparison are described in Section II. The details of PDNGraph construction and PDNNet architecture are presented in Section III. In Section IV, a series of experiments are conducted to comprehensively explain the effectiveness of PDNGraph and PDNNet. The paper is concluded with a summary in Section V.
II Preliminary
II-A IR Drop Analysis
IR drop analysis simulates the voltage drop under the combined effect of parasitic and current flow through the PDN [22]. Generally, PDN is casted as a grid topology network with parameters, that mainly include impedance and conductance.[23, 24, 25, 26, 27, 28, 29, 30, 31] Voltage sources (i.e., power supply rails) and current loads (i.e., cells) respond to simulate the current flowing on the network. The illustration of PDN grid depicts in Figure 2.
Figure 2: The illustration of PDN grid in IR drop analysis.
Hence, the IR drop analysis involves the solution of a large system [32] of equations of the form,
$GV\ =\ J$ (1)
where $G$ is the conductance matrix for the PDN, $J$ is the vector of current sources, and $V$ is the set of voltages at each node in the PDN. The whole procedure [32] iteratively measures voltage at each node in four key consecutive steps as depicted in Figure 3 (a).
  • Step 1: Calculate effective resistance and capacitance between this node, the on-grid voltage sources, and current loads parts.
  • Step 2: Calculate node’s voltage drop induced by global current flows to current loads.
  • Step 3: Acquiring node’s voltage supply transmitted by current flows from voltage source.
  • Step 4: Consider node’s voltage fluctuation influenced by the voltage of neighbor nodes.
The IR drop analysis can thus be formulated as:
$y\ =f\circ g\circ h\circ s(x)$ (2)
where $s$ present voltage, current map and effective resistance calculation, $h$ denotes the voltage drop acquiring, $g$ represents the calculation of voltage source supply, and $f$ denotes the influence from neighbor. Notation $\circ$ represents the connections between two operations. The analysis formulation explicitly indicates that PDN structure and dynamic current pattern variations jointly determine IR drop distribution map on the grid.
(a) Simulation-based IR drop analysis pipeline.
(b) Conventional CNN-based IR drop prediction method.
(c) Proposed PDNGraph construction process.
(d) Proposed PDNNet IR drop prediction framework.
Figure 3: Analysis of simulation-based method, conventional deep-learning prediction methods, our proposed PDNGraph, and PDNNet framework.
Figure 4: Illustration of our PDNGraph construction. Figure better viewed in color mode. From left to right is (a) The whole graph on IC layout. (b) The fine-grained part inside PDN strip. (c) The fine-grained part outside PDN strip.
II-B Problem Formulation
The former IR drop modeling suffers from a high computation burden. Hence, we approach this procedure in a learnable fashion. This task, named IR drop prediction, is of great value in accelerating loops of IC design closure. We identify our learning formulation as follows:
Problem 1. Given a series of data after placement $\{PL\}$, consisting of PDN attributes $P=\{p_{i}\mid 1\leq i\leq N_{p}\}$, standard cells locations $C=\{c_{i}\mid 1\leq i\leq N_{c}\}$ and cell current related features $X=\{x_{i}^{k}\mid 1\leq i\leq N_{c},1\leq k\leq K\}$, our object is to design an algorithm $\mathcal{G}$ to compose these features in a specific manner and propose a learning function $\mathcal{F}$ to intake the processed data to generate an accurate dynamic IR drop distribution map.
II-C Comparison with Other Methods.
Table I presents the difference between PDNNet, Voltus, and other typical models for dynamic IR drop analysis. Voltus is a commercial tool that employs simulation-based method to derive an IR drop map. The results obtained from this tool serve as our ground reference. Within prediction-based methods, XGBoost and CNN represent two commonly employed models in machine learning and deep learning, respectively. Cell-PDN relation refers to the consideration of the physical interconnection between the PDN network and cells. Generality and speed stand as two pivotal parts of the practical usability of these methods. The results are based on our experiments in Section IV.
The PDNGraph data structure and PDNNet model enable our method to generate an accurate IR drop map by mimicking the dynamic current flow that exists in PDN structure, distinguishing it from the abovementioned prediction-based methods. Specifically, PDNGraph uses graph structure to favorably represent the fine-grained PDN structure and intrinsic cell-PDN physical relation. In the meantime, PDNNet leverages a dual-branch GNN-CNN heterogeneous network to achieve a comprehensive perception of both PDN architecture and dynamic IR drop variation. These advancements boost our PDNNet’s performance significantly.
III Method
In this section, we propose a solution pipeline, including a novel graph structure, PDNGraph, and a novel network architecture, PDNNet. The overview of PDNGraph and PDNNet is shown in Figure 3 (c) and Figure 3 (d).
For PDNGraph, it is a directed graph that encompasses the electrical properties of PDN structure and current flow. We assign several current-related features (e.g., leakage power, internal power, switching power) to each node as internal attributes. Then, the inter-node edge is constructed to direct from the high power rail toward the low redistribution metal strip, which follows the actual current flow direction through PDN. As such, PDNGraph is built as a bridge to establish the fine-grained cell-PDN connection.
Furthermore, our proposed PDNNet is a dual-branch heterogeneous aggregation network. Specifically, for the GNN-branch, we employ a graph neural network to aggregate PDNGraph. We adopt different GNN layers in the voltage-drop acquiring and neighborhood influence stages, respectively, which align more closely with the formulation pipeline in Section II-A. As for the CNN-branch, we introduce a 3D encoder and 2D decoder network with several shortcut connections to capture the dynamic IR drop fluctuation. Finally, a small yet effective fusion model merges each branch output to yield the final prediction result. We elaborate on the details of several key designs of our method in the following sub-sections.
III-A PDNGraph
To make appropriate input data for accurate dynamic IR drop prediction, there are two main challenges. Firstly, the current load pattern of cells needs to extract in both the spatial and temporal axis, which is different from other tasks. Secondly, the PDN network characteristics are also critical to impact the voltage drop distribution, which is a noteworthy clue. However, the existing methods stress more on simply leveraging current patterns to approach IR drop prediction, while overlooking the intrinsic cell-PDN physical relation and fine-grained structures of PDN. To counter these limitations, we shed light on characterizing these factors in graph format. We propose a novel graph data structure, named PDNGraph, to comprehensively represent PDN structure and cell-PDN relationship. We elaborate PDNGraph attributes as follows.
Figure 5: Architecture of PDNNet for dynamic IR drop prediction. Our PDNNet consists of two parallel branches: the GNN branch in the upper part and the CNN branch in the lower part. $H,W,T$ denotes the input features’ height, width, and temporal dimension.
III-A1 Graph Construction
We divide the whole layout into uniform grids. The Figure 4 (a) illustrates our PDNGraph outline, where we define the part within a grid as the node of PDNGraph, while the edge connects adjacent nodes in one-hop is restricted only exists in the horizontal and vertical directions. The inter-node distance, corresponding to the resistance distribution, is also assumed uniform across the whole design [18]. In PDNGraph, we opt not to explicitly calculate resistance and capacitance for each edge, for fully believing electrical law can be “learned” by an artificial neural network (e.g., our PDNNet). Hereafter, PDNGraph manifests a consistent manner with the grid network in IR drop analysis.
III-A2 Current Load Attribute
According to the analysis of IR drop modeling, the cells’ current consumption pattern is the key feature of node attributes. Following the protocol of typical work in [18], as power consumption is proportional to current, PDNGraph also opts for cell power as its input feature. The encompassed power features are:
  • Leakage power, also known as static power, is the power dissipated when a small current flowing between its source and drain terminals while in the off state.
  • Internal power, refers to the power consumed by the active components during its operation. Since it is variant along temporal, internal power is a type of dynamic power.
  • Switching power, another type of dynamic power, is the power consumed during the transition of electronic signals switch from one logic level to another (e.g., from 0 to 1).
We compose and assign these features to form node features $x$, $x\in\mathbb{R}{{}^{N\times C}}$, where $N$ denotes the node number in PDNGraph and $C$ represents the input feature dimension.
III-A3 Current Direction Attribute
In the general PDN configuration, the high power supply VDD rails (HPR) transmit current to the low redistribution VDD layer (LRL). Despite taking complete PDN into account, we opt for the most relevant part of IR drop distribution to avoid redundancy. Firstly, without considering the macros, LRL (i.e., M0 or M1) is the layer that directly connects to the cells where most of the IR drop occurs. Secondly, the location where crisscrossed HPR (i.e., M7 in  [33]) directly connects to LRL through vias has the least IR drop violation. To this end, the power supply process can be illustrated that the current flow spread from HPR ends to LRL, along the LRL strip direction towards each node. The voltage drop at this point, as visually explained in Figure 4 (b)(c), also increases gradually when LRL moves away from the HPR. Therefore, we designated the edges’ direction starting from the closest HPR towards the away direction. Since there is no explicit current direction between vertical nodes, we consider them as undirectional edges. Further, we neglect the influence of current on the voltage drop across HPR with respect to their low resistance nature. Notice that our PDNGraph can be adaptive to the uniform or non-uniform PDN with high flexibility, as shown in Figure 4 (a).
III-B PDNNet
III-B1 GNN Branch
Graph Neural Networks (GNNs) have achieved significant success in node analysis tasks (e.g., node classification, node clustering, etc). Instead of stacking the same GNN blocks, here we opt for distinct graph aggregation layers following the formulation pipeline in Section II-A. The architecture is illustrated in the top of PDNNet in Figure 5. In the GNN branch, we first feed the PDNGraph into hidden representations through Multi-Layer Perceptrons ($MLP$). Then, deeper representation is generated by message-passing through voltage-drop acquiring and neighbor influence layers respectively. Finally, the GNN branch representation is exported through the readout layers.
Voltage-Drop Acquiring Layer. To utilize the current-related features to learn the voltage drop, we stack $l$ voltage-drop acquiring blocks to form this layer. To better acquire one nodei’s voltage drop representation, we sum all the neighbors’ voltage drop features,
$m_{i}\ ={\displaystyle\sum_{k\in n(i)}\left(v_{k}^{(l-1)}\right)},$ (3)
$n$ is nodei’s neighbors and $m_{i}$ is the neighbors’ feature summation. Therefore, we update the feature of this node by concatenating all neighbor effects,
$v_{i}^{(l)}=f_{\phi}(v_{i}^{(l-1)}\ \|\ m_{i}),$ (4)
[…]
Neighbor Influence Layer. Since the node features have already captured the characteristics of voltage drop through the above layers, the objective of this layer is to learn and refine the influence of neighbors on each node. We therefore update the PDNGraph to an undirected graph while still preserving its topological structure. Subsequently, learnable function blocks $f_{\theta}$ are employed to update the node representation using both the original node features and the ones from its neighbors,
$v_{i}^{(l)}=f_{\theta}(v_{i}^{(l-1)}\ \|\ v_{k}^{(l-1)})\ \ k\in n(i),$ (5)
where $\theta$ is a set of learnable parameters of graph convolution layers.
Voltage Representation Readout. After the preceding two layers of message-passing, we employ an $MLP$ layer followed by a $tanh$ activation to yield the PDN-related representation.
$y^{gnn}=tanh(MLP(x^{(l)})),$ (6)
III-B2 CNN Branch
CNN branch in PDNNet is performed as a parallel temporal branch of GNN’s. Although GNN-branch can obtain considerable performance, CNN branch is indispensable for perceiving current alteration along the time axis in dynamic IR drop prediction. Details can be founded in the experiment Section IV. To achieve temporal perception, we directly adopt a classical encoder-decoder structure with 2 shortcut connections. The architecture is present in the bottom of Figure 5. More advanced architectures can be applied here, but this is not the focus of this work. Details of the CNN branch are described as follows. The intake features share the same as those of the GNN branch. We map the graph node features onto a 2D space and insert a dummy dimension instead of the original channel. The transformed feature shape is $x\in\mathbb{R}{{}^{H\times W\times T\times C}}$, where $T$ denotes the number of input features and $C$ means feature channels through the representation process. The input feature is first fed into a 3D encoder network that goes through $l$ encoder layers, each layer comprise of two 3D convolutional and down-sampling blocks. After feature encoding, the output embedding shape converts to $x^{cnn}\in\mathbb{R}{{}^{\frac{H}{2l}\times\frac{W}{2l}\times\frac{T}{l}\times C}}$, where $l$ opt as 3 in our model. At this point, we shrink the temporal channels since they have been sufficiently extracted.
$x^{cnn}\in\mathbb{R}{{}^{\frac{H}{2l}\times\frac{W}{2l}\times\frac{T}{l}\times C}}\ \longrightarrow\ x^{cnn}\in\mathbb{R}{{}^{\frac{H}{2l}\times\frac{W}{2l}\times C}}$ (7)
Subsequently, comprising $l$ layers of deconvolution and up-sampling blocks, the 2D decoder network maintains mirror symmetry in spatial shape with the 3D encoder counterpart at each level. To further assist the recovery of low-level information from the original feature, we fuse the 4D features, as shown in Figure 5, from the left side into each layer’s 2D block. Finally, the CNN network outputs IR drop temporal representation $y^{cnn}\in\mathbb{R}{{}^{H\times W\times 1}}$.
TABLE II: Evaluation results on CircuitNet dataset. Our PDNNet achieves consistent performance gain on five evaluation metrics in both large and small datasets. Red denotes percentage of improvement. $\downarrow$ means “lower is better”, $\uparrow$ means “higher is better”.
Method CircuitNet [33]
[…]
PowerNet [18] 0.107 0.59 0.74 0.72 0.62
PDNNet (Ours) 0.031(-71.0%) 0.90(+52.5%) 0.95(+28%) 0.75(+4.2%) 0.63(+1.6%)
MAVIREC [19] 0.038 0.47 0.89 0.74 0.62
PDNNet (Ours) 0.031(-18.4%) 0.90(+91.5%) 0.95(+6.7%) 0.75(+1.4%) 0.63(+1.6%)
III-C Fusion Module
In the last part, we fuse upstream embedding together to get a comprehensive representation. We first align the feature dimensions by flattening the temporal representation $y^{cnn}\in\mathbb{R}{{}^{H\times W\times 1}}$ into $y^{cnn}\in\mathbb{R}{{}^{HW}}$. Then, a concatenation function is imposed to merge these features and pass them through an $MLP$ layer as a regressor to generate the final IR drop prediction.
$z^{prediction}\ =\ MLP(y^{gnn}\ \ \|\ \ y^{cnn}),$ (8)
III-D Discussion
We notice that some literature attempts to involve PDN features. The previous work [34] targets static IR drop prediction that regards the PDN as a 2D density map, where each pixel denotes the total number of PDN strips covering this location. However, our PDNNet manifests an obvious difference with this method. First, our method targets more difficult dynamic IR drop prediction. Second, we construct a graph structure based on the intuition of IR drop analysis to finely represent the cell-PDN physical relation, compared to the image-like PDN density map. Third and finally, we design a dual-branch GNN-CNN heterogeneous network to favorably aggregate PDN attributes and temporal variation for accurate dynamic IR drop prediction, compared to vanilla image encoder-decoder network (like Figure 3 (b)).
IV Experiments
In this section, we first conduct experiments on large public datasets to evaluate the performance of our model. Then, we perform extra experimental ablation studies to explore the performance gained from several key designs.
IV-A Experimental Setting
Environment Setups. We implement our method using PyTorch and the DGL graph learning framework. The experiments are conducted on a Linux machine with an NVIDIA A800 GPU, two 32-core Intel Xeon Processors (Ice Lake), and 1024GB memory.
TABLE III: The statistics of Mini-CircuitNet. Samples randomly sample from the origin superset. Note that training & testing datasets have no intersection in designs.
{tabu}
l—c—c—ccc Design Number of Samples Netlist Statistics
#Cells #Nets Cell Area ($\mu m^{2}$)
Train RISCY-a 20 44,836 80,287 65,739
RISCY-b 22 30,207 58,452 69,779
RISCY-FPU-a 21 61,677 106,429 75,985
RISCY-FPU-b 17 47,130 84,676 80,030
Test zero-riscy-a 13 35,017 67,472 58,631
zero-riscy-b 7 20,350 45,599 62,648
Datasets. CircuitNet [33, 35] is the large-scale public dataset in development that provides IC design used in the real industrial scenario. Unlike most previous work using small datasets for validation, CircuitNet-N28 (28nm version) provides more than 10,000 samples from 6 basic RTL designs (54 different synthesized netlists) with a wide variety of macros (3 types), frequencies (50/200/500 MHz), and back-end flow (variant in settings of utilization, macro placement, power mesh, filler insertion, etc), which is favorably appropriate to examine the actual performance of dynamic IR drop prediction models. We follow the same protocol as CircuitNet [33], to split the training and testing sets. That is, 4 designs (RISCY-a, RISCY-b, RISCY-FPU-a, RISCY-FPU-b) with 7078 samples are for training, and the rest 2 designs (zero-riscy-a and zero-riscy-b) with 3164 samples are for testing. CircuitNet-N28 leverages Innovus and Voltus flow to obtain vectorless dynamic IR drop solutions. In our experiment, we extract the PDN features from the Design Exchange Format (DEF) files while incorporating the power information that was initially provided to complete our PDNGraph construction. Besides, to avoid unnecessary training burden, we also collect a subset of CircuitNet, named Mini-CircuitNet, for conducting the ablation studies in Section IV-C. To do so, we sample 100 total samples from the original CircuitNet dataset, where 80 and 20 samples are for the training and testing phases respectively. As the sampling operation is random, we opt not to deliberately pursue a uniform number of samples in each design. Notice that the above training and validation datasets have no intersection across designs. Refer to Section IV-A for more details.
Implementation Details. We train all models, including baseline methods, using Adam optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, the initial learning rate is set as 0.0008 and apply a low weight decay of $0.0001$. We gradually reduced the initial learning rate with the cosine annealing strategy. Datasets pre-processing follows [33]. Note that we do not impose any data augmentation transform during the training stage.
TABLE IV: Evaluation results on Mini-CircuitNet dataset. Red denotes percentage of improvement. $\downarrow$ means “lower is better”, $\uparrow$ means “higher is better”.
Method Mini-CircuitNet
[…]
PowerNet [18] 0.171 0.62 0.75 0.70 0.60
PDNNet (Ours) 0.034(-80.1%) 0.87(+40.3%) 0.94(+25.3%) 0.75(+7.1%) 0.62(+3.3%)
MAVIREC [19] 0.056 0.82 0.92 0.71 0.58
PDNNet (Ours) 0.034(-39.3%) 0.87(+6.09%) 0.94(+2.2%) 0.75(+5.6%) 0.62(+6.9%)
IV-B Main Experiment
We first compare our PDNNet with the state-of-the-art CNNs from the literature.  The first baseline for comparison is PowerNet [18], which is the first paper to perform dynamic IR drop prediction with deep-learning method.  The second is MAVIREC [19], a UNet-like model learning on dynamic IR drop with input vector, which is currently the state-of-the-art among the vector-based or vector-less methods. In general, the CircuitNet dataset expects neural networks to have strong generalization and representation capabilities due to its diverse design compositions. Thus, for evaluation metrics, we report both the accuracy metric (i.e., Normalized Mean Absolute Error (NMAE)), coefficient of determination (i.e., R-square ($R^{2}$)) and the graph evaluation criterion (i.e., Pearson, Spearman, Kendall correlation coefficient), to comprehensively assess the models’ performance. We abbreviate Pearson to Pear for shorten, Spearman and Kendall convert to Spea and Kend as the same. We calculate NMAE metric as $\frac{1}{N}\sum_{i=0}^{{N}-1}\frac{{\left|y_{i}-\tilde{y}_{i}\right|}}{max(y_{i})-min(y_{i})}$.
Accuracy Comparison. The Table IV shows the results. Compared to the PowerNet, PDNNet achieves over 71% improvement (0.031 vs. 0.107) in NMAE metrics, which depicts the incomparable generalization of our PDNNet, other metrics also support this viewpoint. The weak performance of PowerNet is primarily due to its simplistic network design and non-mainstream 3D learning strategy, which limits its ability to learn deeper features related to IR drop. The typical CNN architecture, MAVIREC, still has a great margin compared with our PDNNet (0.038 vs. 0.031 in NMAE, 0.47 vs. 0.90 in $R^{2}$, 0.89 vs. 0.95 in Pearsonr, etc). It is worth noting that even in the experiments using Mini-CircuitNet with fewer samples (rightmost column of Table IV), our PDNNet has consistent and significant performance gains over all the five metrics, which implicitly demonstrates the limitation of overlooking PDN attributes during IR drop representation. Accuracy-computation comparison is illustrated in Figure 7. Visualization comparison results are shown in Figure 6. The figures from the left to right columns are macro positions in the floorplan, results from MAVIREC , results from our PDNNet, and ground truth label. The comparison results clearly show that our PDNNet generates IR drop prediction map with high structural fidelity, which demonstrates the effectiveness of our method.
AUC Comparison. We also calculate the Area Under the Curve (AUC) score to measure model representation performance, PDNNet achieves 0.9748, outperforming the result of 0.9599 in MAVIREC.
Case 1
Refer to caption Refer to caption Refer to caption Refer to caption
Case 2
Refer to caption Refer to caption Refer to caption Refer to caption
Case 3
Refer to caption Refer to caption Refer to caption Refer to caption
Floorplan  MAVIREC [19] Our PDNNet   Ground Truth
Figure 6: Qualitative evaluation of CircuitNet dataset. The green polygons in the floorplan layout represent macros. Our PDNNet generates IR drop prediction map with high structural fidelity, which demonstrates the effectiveness of our method.
Accuracy-computation trade-off. Accuracy and computation cost are two crucial criteria for comprehensive model evaluation. The whole accuracy-computation trade-off of different methods is illustrated in Figure 7. We compare the accuracy and inference time between commercial tools, other compared methods, PDNNet variants, and our PDNNet. The inference time of our PDNNet is slightly higher than the MAVIREC (0.088s vs. 0.065s), while significantly lower than PowerNet and commercial tool Voltus (0.088s vs. 52.1s/60-120s). PDNNet and MAVIREC have similar inference costs as they both adopt 3D-CNN structures to aggregate temporal dimensions. For PowerNet, the unique max-CNN structure brings a large computation burden that costs much more inference time. The commercial tool, Voltus, uses a simulation-based method to acquire an IR drop distribution map and adopt it as our ground reference. Overall, the average runtime of PDNNet is around 0.11s (0.088s for model inference, 0.016s for data preprocessing), which performs over 545$\times$ speedup compared to commercial tool (1-2 minutes approximately).
IV-C Ablation Studies
In this section, we investigate the effectiveness of each individual technique by ablation studies. We perform two series of experiments to investigate the following research questions: RQ1. How effective is the mixure of different message-passing layers in the GNN branch compared with each individual of them? RQ2. How effective is the dual-branch heterogeneous PDNNet compared with the pure CNN or GNN methods?
Figure 7: Accuracy-computation trade-off on Mini-CircuitNet for different methods in the inference phase. The horizontal axis shows the inference cost (i.e., whole inference time or model inference time) per sample. Whole inference time includes data processing and model inference.
Physical Intuitive GNN Branch Design (RQ1). We perform a preliminary exploration of GNN branch design by constructing two variants, a full layers of voltage-drop acquiring blocks and a full layers of neighbor influence blocks. We train the above models and PDNNet on Mini-CircuitNet with the same settings. Results in Table V, show that two variants have similar performance. The voltage-drop acquiring blocks is slightly higher than its neighbor influence counterpart (0.0348 vs. 0.0346) under the NMAE metric while exceeding in $R^{2}$ (0.8668 vs. 0.8665). However, by merging these distinctive layers in a specific manner, we observe a performance improvement in our PDNNet’s GNN branch (-2.9% and +0.3% in NMAE and $R^{2}$ respectively). These results reinforce the rationality of the formulation in Section II-A that shallow layers of GNN branch utilize current-related features to learn the preliminary representations of voltage drop, while deeper layers is to refine nodes’ voltage drop features influenced by their neighbors.
Dual-Branch Heterogeneous Architecture Analysis (RQ2). We perform a controlled study to assess PDNNet’s performance versus each variant model’s. The model set includes single-branch CNN and dual-branch CNN, plus single-branch GNN and dual-branch GNN. The dual-branch variants stand for using the same architecture as PDNNet while simply substituting with two identical branches (CNN or GNN). Table VI contains the comparison of the results, while the accuracy-speed comparison is shown in the inner box of  Figure 7.
TABLE V: Results of constructing GNN branch with different message-passing layers.
[…]
Full Voltage-Drop Acquiring Layers 0.0348 0.8668
Full Neighbor Influence Layers 0.0346 0.8665
Mixed GNN Layers (our GNN branch) 0.0336 0.8695
TABLE VI: Influence of GNN & CNN branch of PDNNet.
[…]
CNN 0.0447 0.8253
0.0361 0.8602
GNN 0.0420 0.8332
0.0347 0.8659
PDNNet 0.0336 0.8695
We have the following observations. First, although CNNs dominate the IR drop prediction in previous works, we interestingly observe that our GNN branch’s performance and inference speed both match or exceed its CNN-based counterpart, whilst with a simple PDN perception and relatively shallow layers (two GNN layers, each accompanied with one GNN block). This phenomenon is largely due to our specifically designed PDNGraph data structure which effectively mimics current flow and fine-grained cell-PDN relation within the PDN grid. Second, compared to a single-branch, dual-branch architecture indeed boosts the prediction performance (19.2% gains in CNN, 17.4% gains in GNN, under NMAE). However, both of these variants do not outperform our heterogeneous PDNNet (0.0361/0.0347 vs. 0.0336, and 0.8602/0.8659 vs. 0.8695), which demonstrate the PDN configuration and dynamic current pattern representations are two complementary aspects for effective IR drop prediction.
V Conclusion
This paper introduces a novel approach to accurately predict dynamic IR drop on the power delivery network in IC design. We present a novel graph structure, PDNGraph, to properly represent the PDN configuration and the fine-grained cell-PDN relation. Moreover, we propose a novel network, PDNNet, a heterogeneous network with GNN-CNN branches to favorably perceive the above crucial features. Extensive experiments on a large public dataset demonstrate that our PDNNet outperforms the existing state-of-the-art CNN-based methods by a large margin (up to 39.3% in NMAE) and achieves a 545$\times$ speedup compared to the commercial tool with minimal error. We hope our work can foster further inspiration for EDA researchers.